in a g.p series 2n number is present.S1 is the sum of numbers occupying even position and S2 is sum of odd numbers.find common ratio.

Grade:

1 Answers

Arun
5 years ago
Dear student
 
Let G.P. is
a, ar, ar^2, ar^3, …....
Now
S1 = ar (r^n  – 1) / (r – 1)
and
S2 = a (r^n  – 1) / (r – 1)
hence
s1 / s2 = r